OpenStudy (anonymous):

solve by elimination 1. x-y+3z=8 2.3x+2y+=5 3.4x+y+6z=11

OpenStudy (anonymous):

look at number 2 ... is that correct ?

OpenStudy (anonymous):

i mean have u typed it rightly?

OpenStudy (anonymous):

3x+2y\( \color{red}{+} \)=5

OpenStudy (anonymous):

3x+2y+5z=5

OpenStudy (anonymous):

well i solved it...first look at these: x-y+3z=8 4x+y+6z=11 so we can write: 5x + 9z = 19 \(\Large \color{red}{ (1)} \)

OpenStudy (anonymous):

and now look at these two equations too: 3x+2y+5z=5 4x+y+6z=11 so we can write the second one in this way: -8x - 2y - 12z = -22 and we had 3x+2y+5z = 5 too so it would be: -5x -7z = -17 \( \Large \color{red}{(2)} \)

OpenStudy (anonymous):

and can u find x and z from the equations 1 and 2 ?

OpenStudy (anonymous):

that explanation made sense ?

OpenStudy (anonymous):

yes thank you @PFEH.1999

OpenStudy (anonymous):

you're welcome ;)